Enter your answer in the provided box. In the Mon
Answer #1

Answer – We are given, reaction – Ni(s) + 4CO(g) <----> Ni(CO)4(g)

ΔGof Ni(CO)4(g) = -587.4 kJ/mol , ΔGof CO(g) = -137.3 kJ/mol , T = 76+273 = 349 K

Now we need to calculate the ΔGorxn

We know formula

ΔGorxn = sum of the ΔGof product - sum of the ΔGof reactant

So, ΔGorxn = [ΔGof Ni(CO)4(g)] – [4* ΔGof CO(g)]

ΔGorxn = (-587.4 kJ/mol) – ( 4*-137.3 kJ/mol)

           = (-587.4 kJ/mol) – (-549.2)

           = -38.2 kJ/mol

            = -3.82*104 J/mol

Now we know the formula for calculating the equilibrium constant form the standard free energy.

ΔGorxn = -RT*lnK

-3.82*104 J/mol = - 8.314 J/mol.K * 349 K * ln K

ln K = -3.82*104 J/mol / - 8.314 J/mol.K * 349 K

       = 13.16

So, taking anitln form both side

K = 5.19*105